3x2+7x+2=0
3x2-16x+5=0
7x2-9x+2=0
10u2+17u+3=0

Answer by checkley75(3666)

3X^2+7X+2=0
(3X+1)(X+2)=0
3X+1=0
3X=-1
X=-1/3 ANSWER.
X+2=0
X=-2 ANSWER.
------------------------------------------
3X^2-16X+5=0
(3X-1)(X-5)=0
3X-1=0
3X=1
X=1/3 ANSWER.
X-5=0
X=5 ANSWER
-----------------------------------------
7X^2-9X+2=0
(7X-2)(X-1)=0
7X-2=0
7X=2
X=2/7 ANSWER.
X-1=0
X=1 ANSWER.
------------------------------------------
10U^2+17U+3=0
(5U+1)(2U+3)=0
5U+1=0
5U=-1
U=-1/5 ANSWER.
2U+3=0
2U=-3
U=-3/2 ANSWER.